# LLM-RankFusion: Mitigating Intrinsic Inconsistency in LLM-based Ranking

### v1

@@ Abstract
Ranking passages by prompting a large language model (LLM) can achieve promising performance in modern information retrieval (IR) systems. A common approach is to sort the ranking list by prompting LLMs for pairwise comparison. However, sorting-based methods require consistent comparisons to correctly sort the passages, which we show that LLMs often violate. We identify two kinds of intrinsic inconsistency in LLM-based pairwise comparisons: order inconsistency which leads to conflicting results when switching the passage order, and transitive inconsistency which leads to non-transitive triads among all preference pairs. In this paper, we propose LLM-RankFusion, an LLM-based ranking framework that mitigates these inconsistencies and produces a robust ranking list. LLM-RankFusion mitigates order inconsistency using in-context learning (ICL) to demonstrate order-agnostic comparisons and calibration to estimate the underlying preference probability between two passages. We then address transitive inconsistency by aggregating the ranking results from multiple rankers. In our experiments, we empirically show that LLM-RankFusion can significantly reduce inconsistent pairwise comparison results, and improve the ranking quality by making the final ranking list more robust. Our code is publicly available at https://github.com/XHMY/LLM-RankFusion
**footnotetext: Equal Contribution.footnotetext: Corresponding Author.
@@ 1 Introduction
Large language models (LLMs) have demonstrated strong zero-shot and few-shot capabilities in many natural language processing tasks [1, 46, 47]. This enables the effective integration of these LLMs in modern information retrieval (IR) systems [45, 20]. Without supervised training on labeled data in a specific task, LLMs can adapt to the task by prompt and pipeline design. Recent work has tried to apply LLMs in text ranking and shown promising performance [32, 5, 38, 28]. Text ranking is an important task in modern recommendation systems and search engines, which refines the order of the retrieved passages to improve the output quality [25]. Traditional supervised text ranking methods mostly relied on large amounts of human-annotated labels [2, 3, 42, 29]. Typical LLMs-based ranking approaches prompt LLMs to generate partial orders, comparisons, or relevance scores without additional training in advance.
@@
Despite the great potential of LLMs in passage ranking, they can also suffer from significant inconsistency. Previous work shows LLM-based ranking is sensitive to the order of input passages in the prompt, which stems from the positional bias of LLMs [44, 27, 39]. We identified this as order inconsistency. In a pairwise comparison between two passages, the results can conflict before and after swapping the passages. This makes ranking framework hard to merge the comparison results into a correct ranking list. On the other hand, even if we can fully mitigate the order inconsistency within a single comparison, there is still a concern about inconsistency among different comparisons. For example, PRP-Sort [32] uses sorting algorithms to efficiently produce a full ranked list from pairwise comparisons. However, these sorting algorithms typically assume the transitivity of comparisons to correctly sort the passage, which we show that LLMs often cannot guarantee. (e.g. $d_{1}\succ d_{2},d_{2}\succ d_{3}\Rightarrow d_{1}\succ d_{3}$, where $d_{i}$ represents a passage and $\succ$ means "preferred to"). We identify this as transitive inconsistency, which has been ignored in previous works. Due to the transitive inconsistency, we show that the ranker’s performance is highly sensitive to the initial input order of the retrieved documents. Ranking systems expect to produce a robust ranking list to present to users, but if different initial orderings can lead to significant variant ranked lists, the reliability of the LLM’s rankings can be brought into question.
We propose the LLM-RankFusion framework as shown in Figure 1 to produce a consistent ranking list by mitigating the above inconsistencies. To mitigate order inconsistency, we first use calibration to address the conflict before and after swapping the passages. It calculates the preference probability based on the logit output, which then produces the preference without bias in position. To further let LLM realize the preference should be order agnostic, we propose the in-context learning (ICL) ranking prompt. It uses an example to demonstrate the swapping of passages doesn’t affect the preference. While we can greatly reduce order inconsistency by ICL and calibration, the improved pairwise comparison still does not address transitive inconsistency directly. We use rank aggregation to further mitigate the transitive inconsistency. Rank aggregation is a commonly used post-process method in combining multiple ranking lists and yielding a robust ranking result [12, 23, 36]. We aggregate the ranking list proposals from LLM-based rankers with different underlying sorting algorithms or different LLMs that are responsible to make preference decisions.
@@
  • We propose to use ICL and calibration to address order inconsistency. The improvement is significant in most LLMs.
  • We bridge the area of the LLM-based ranking with rank aggregation for robustness and mitigating the impact of transitive inconsistency.
@@ 2 Related Work
LLM-based ranking approaches
have been developed with distinct ranking schemes. Pointwise approaches [21, 35, 10] aim to estimate the relevance between a query and a single document. Listwise [38, 28] ranking methods aim to directly rank a list of documents by inserting the query and document list into an LLM’s prompt and instructing it to output the reranked document identifiers, though they rely on the strong capability of LLMs, suffer from positional bias and are sensitive to document order in the prompt [49]. Pairwise ranking methods [32] provide the query and a pair of documents to the LLM, which is instructed to generate the identifier of the more relevant document; these pairwise comparisons are then aggregated using efficient sorting algorithms like Heapsort or Bubblesort to produce the final ranking. The Setwise approach [50] is also proposed to compare a set of documents at a time to further improve efficiency.
Rank aggregation
has been widely used in many information retrieval tasks [14, 12]. Previous works [31, 15] also use aggregation to form the ranking list from pairwise comparison results. We employ the Borda count method [9] to aggregate different ranking lists in our paper. Borda count assigns a score to each item based on its position in each input ranking and sums these scores to produce the final aggregate ranking. While other rank aggregation methods exist, such as Markov Chain-based methods [12], supervised learning approaches [26], and unsupervised methods like Kemeny rank aggregation [18], many of these are either NP-hard or not specifically designed for list aggregation. By using the simple and computationally efficient Borda count method, we aim to demonstrate the power of combining LLM-based ranking with aggregation, even without resorting to more complex aggregation techniques.
3 Inconsistency Concerns for LLM-based ranking
3.1 Inconsistency of pairwise comparisons
  1. 1.
  2. 2.
Under the pairwise ranking approach, each LLM query produces a pairwise comparison results on $d_{i},d_{j}$, the result can be $d_{i}\succ d_{j}$ or $d_{j}\succ d_{i}$, where $d$ represents a passage and $\succ$ means "preferred to". While we focus on pairwise comparisons in this work, the definition of these two types of inconsistency can also be extended to listwise or other LLM-based ranking schemes. We follow PRP-Sort [32] to handle order inconsistency as a tie. But we know that not all these ties represent equal relevance passages given the ground truth relevance labels.
3.2 Inconsistency Measurement
@@
A tournament graph is a complete graph with directed edges representing preference between two nodes and undirected edges representing ties. We use a tournament graph to study the transitive inconsistency of LLM-based ranking. We construct the tournament graph from the pairwise comparison results by setting all order inconsistent pairs to ties. We follow [19] and use the number of inconsistent triads to measure the transitive inconsistency. As shown in Table 2, the level of inconsistency can vary depending on the LLM. We find that LLMs with larger size typically suffer from less transitive inconsistency.
@@
Table 2: The number of inconsistent triads in the tournament graph. A higher number of inconsistent triads indicates more inconsistency. Type-1 triads refer to $d_{i}=d_{j},d_{j}=d_{k},d_{k}\succ d_{i}$. Type-2 triads refer to $d_{i}=d_{j},d_{i}\succ d_{k},d_{k}\succ d_{j}$. The total inconsistent triads refer to the sum of all kinds of inconsistent triads. The maximum number of inconsistent triads for 100 passages is 161684 [19]. This tournament graph is constructed from the comparisons without ICL and calibration.
@@ 3.3 The impact of inconsistency
BM25 "Hard List" BM25 "Hard List"
@@
Table 3: NDCG@10 of ranking from different initial orders. We demonstrate the effects of inconsistency shown in Figure 2 by constructing a hard initial ordering for an LLM-based ranker based on Bubblesort. This "hard list" is the inverse order of the ranking result obtained by performing the PRP-Allpair ranking scheme [32]. Note that the Heapsort results are largely unaffected by it, suggesting that this specific "hard list" is for Bubblesort only.
As shown in Tables 1 and 2, LLM-based rankers can exhibit significant amounts of inconsistency across judgments. As illustrated in Figure 2, applying sorting-based ranking schemes on the basis of non-transitive pairwise comparisons can produce volatile result rankings that are highly sensitive to the initial order of candidate passages. This can have particularly adverse effects if that initial ordering is a "hard list" for the chosen sorting method, as demonstrated in Table 3. A hard list is an initial order of passages where the high-relevance passages require many comparisons to be moved to the front of the ranking, increasing the likelihood of encountering a transitive inconsistency which blocks the promotion of the passage. A hard list for one sorting algorithm may not be as hard of a list for another, which is demonstrated in Table 3. By aggregating the full ranked lists from multiple sorting algorithms, we can mitigate the worst-case effects of inconsistency while producing more robust final rankings.
@@ 4.1.1 In-Context Learning (ICL)
We design the ICL prompt to utilize the few-shot ability [4] of LLMs to mitigate order inconsistency. The prompt provides the LLM with an example pairwise comparison for both order permutations as shown in Figure 1. This demonstration illustrates that the task is to compare the passages based on their relevance to the query and not selecting the passage based on its position in the prompt. As shown in Table 1, using ICL can balance the probability of the LLM selecting a passage from either position.
@@ 4.2 LLM Ranking Aggregation
In a passage ranking task, we can generate multiple proposed ranking lists using different ranking settings. These settings can include different sorting algorithms to build fully ranked lists from pairwise comparisons, as well as other factors like a specific LLM used for the preference query. Given the inconsistency discussed in 3, we know that any configuration of these settings can result in noisy rank results. We cannot assume we know a priori which proposed rank settings is the best, so choosing a single setting that will be affected the least by inconsistency is difficult.
@@
5:     $L_{i}\leftarrow$ PairwiseRanking($q$, $D$, $R_{i}$) $\triangleright$ Generate rank list using rank setting $R_{i}$
@@ 4.2.2 Aggregation Across LLMs
5 Experiments
5.1 Experimental Setup
We utilize test sets from TREC, a standard dataset for information retrieval research. Specifically, we use the top 100 passages retrieved by BM25 [22, 34] for each of the 43 queries associated with the 2019 test set TREC-DL2019 [8]. Our results are based on the re-ranking of these 100 passages. The LLM ranking scheme is implemented under the same experimental setting of PRP-Sort [32].
We evaluate our results using two metrics: Normalized Discounted Cumulative Gain (NDCG) and average Kendall tau Distance. NDCG is a standard metric used to evaluate the quality of ranked retrieval results. It accounts for the position of relevant documents in the ranking, assigning higher importance to documents appearing earlier.
Kendall tau distance is a metric used to measure the dissimilarity between two rankings. For any specific group of ranked list proposals, the volatility can be measured as the average Kendall-tau distance between any two ranked lists constructed from different initial orderings. If this average distance is high, it implies high inconsistency; the final ranking is very sensitive to initial ordering. To evaluate the average Kendall-tau distance for our aggregation pipeline, we construct $n$ ranking lists from $n$ different initial orders. This step is only for evaluation purposes and is not necessary in LLM-RankFusion. The $KT_{avg}\in[0,1]$ is defined as
$KT_{avg}=\frac{1}{|Q|}\sum_{i=1}^{|Q|}\frac{1}{{n\choose 2}}\sum_{p=1}^{n}\sum_{q=p+1}^{n}KT^{pq}_{i}$ (6)
where $|Q|$ is the number of queries, $n$ is the number of unique initial orderings, and $KT^{pq}_{i}$ is the Kendall-tau distance between two rankings for query $i$, $R^{p}_{i}$ and $R^{q}_{i}$, starting from initial orders $p$ and $q$.
@@ 5.1 Experimental Setup
Table 4: NDCG@10 results of PRP-Sorting with Bubblesort, starting from the BM25 initial order. The experiment shows an ablation study of using in-context learning and calibration to improve ranking performance. These two methods aim to address order inconsistency.
@@ 5.2 Addressing Order Inconsistency
We have shown that in-context learning can help to balance the average probability of the two choices given during pairwise ranking in Table 1. In Table 4, we can see that solely using ICL can help improve the ranking performance in most LLMs. Some encoder-decoder models like Flan-UL2 may intrinsically suffer from less positional bias than other decoder-only LLMs [24], so we don’t expect much improvement with using ICL for those models. The calibration addresses order inconsistency by calculating preference probability based on comparison from both positions. The improvement from solely using the calibration is also significant, as shown in Table 4. We also find that using these two methods individually cannot improve ranking performance for Flan-UL2, but combining them can bring improvement compared to the baseline.

### v2

@@ Abstract
Ranking passages by prompting a large language model (LLM) can achieve promising performance in modern information retrieval (IR) systems. A common approach to sort the ranking list is by prompting LLMs for a pairwise or setwise comparison which often relies on sorting algorithms. However, sorting-based methods require consistent comparisons to correctly sort the passages, which we show that LLMs often violate. We identify two kinds of intrinsic inconsistency in LLM-based pairwise comparisons: order inconsistency which leads to conflicting results when switching the passage order, and transitive inconsistency which leads to non-transitive triads among all preference pairs. Our study of these inconsistencies is relevant for understanding and improving the stability of any ranking scheme based on relative preferences. In this paper, we propose LLM-RankFusion, an LLM-based ranking framework that mitigates these inconsistencies and produces a robust ranking list. LLM-RankFusion mitigates order inconsistency using in-context learning (ICL) to demonstrate order-agnostic comparisons and calibration to estimate the underlying preference probability between two passages. We then address transitive inconsistency by aggregating the ranking results from multiple rankers. In our experiments, we empirically show that LLM-RankFusion can significantly reduce inconsistent comparison results, improving the ranking quality by making the final ranking list more robust. Our code is available at https://github.com/XHMY/LLM-RankFusion
**footnotetext: Equal Contribution.
@@ 1 Introduction
Large language models (LLMs) have demonstrated strong zero-shot and few-shot capabilities in many natural language processing tasks [1, 46, 47]. This enables the effective integration of these LLMs in modern information retrieval (IR) systems [45, 20]. Without supervised training on labeled data in a specific task, LLMs can adapt to the task by prompt and pipeline design. Recent work has tried to apply LLMs in text ranking and shown promising performance [32, 5, 38, 28]. Text ranking is an important task in modern recommendation systems and search engines, which refines the order of the retrieved passages to improve the output quality [25]. Traditional supervised text ranking methods mostly rely on large amounts of human-annotated labels [2, 3, 42, 29]. Typical LLM-based ranking approaches prompt LLMs to generate partial orders, comparisons, or relevance scores without additional training in advance.
@@
Despite the great potential of LLMs in passage ranking, they can also suffer from significant inconsistency. Previous work shows LLM-based ranking is sensitive to the order of input passages in the prompt, which stems from the positional bias of LLMs [44, 27, 39]. In a pairwise comparison between two passages, the results can conflict before and after swapping the passages. We identify this as order inconsistency. Even if we can fully mitigate the order inconsistency within a single comparison, there is still a concern about inconsistency among multiple different comparisons. For example, PRP-Sort [32] uses sorting algorithms to efficiently produce a full ranked list from pairwise comparisons. However, these sorting algorithms typically assume the transitivity of comparisons to correctly sort the passage, which we show that LLMs often cannot guarantee. (e.g. $d_{1}\succ d_{2},d_{2}\succ d_{3}\Rightarrow d_{1}\succ d_{3}$, where $d_{i}$ represents a passage and $\succ$ means "preferred to"). We identify this as transitive inconsistency, which has been ignored in previous works. Due to the transitive inconsistency, we show that the ranker’s performance is highly sensitive to the initial input order of the retrieved documents. Ranking systems expect to produce a robust ranking list to present to users, but if different initial orderings can lead to significant variant ranked lists, the reliability of the LLM’s rankings can be brought into question. Our analysis of this inconsistency can be easily extended to setwise and listwise comparisons, as both ranking schemes rely on relative preferences and are fundamentally pairwise in nature. This broad applicability highlights that identifying and addressing these inconsistencies in pairwise ranking is relevant for improving the reliability of any ranking scheme based on relative preferences. We propose the LLM-RankFusion framework as shown in Figure 1 to produce a consistent ranking list by mitigating the above inconsistencies. To mitigate order inconsistency, we first use calibration to address the conflict before and after swapping the passages. It calculates the preference probability based on the logit output, which then produces the preference without bias in position. To further let LLM realize the preference should be order agnostic, we propose the in-context learning (ICL) ranking prompt. The ICL prompt uses an example to demonstrate the swapping of passages doesn’t affect the preference. While we can greatly reduce order inconsistency by ICL and calibration, the improved pairwise comparison still does not address transitive inconsistency directly. We use rank aggregation to further mitigate the transitive inconsistency. Rank aggregation is a commonly used post-process method in combining multiple ranking lists and yielding a robust ranking result [12, 23, 36]. We aggregate the ranking list proposals from LLM-based pairwise and setwise rankers with different underlying sorting algorithms or different LLMs that are responsible to make preference decisions.
@@
  • We address order inconsistency by ICL and calibration. The improvement is significant in most LLMs.
  • We bridge the area of LLM-based ranking with rank aggregation to mitigate the impact of transitive inconsistency in pairwise and setwise comparisons.
@@ 2 Related Work
Table 2: The number of inconsistent triads in the tournament graph. A higher number indicates more inconsistency. Type-1 triads refer to $d_{i}=d_{j},d_{j}=d_{k},d_{k}\succ d_{i}$. Type-2 triads refer to $d_{i}=d_{j},d_{i}\succ d_{k},d_{k}\succ d_{j}$. The total inconsistent triads refer to the sum of all kinds of inconsistent triads. The maximum number of inconsistent triads for 100 passages is 161684 [19]. This tournament graph is constructed from the comparisons without ICL and calibration.
LLM-based ranking approaches have been developed with distinct ranking schemes. Pointwise approaches [21, 35, 10] aim to estimate the relevance between a query and a single document. Listwise [38, 28] ranking methods aim to directly rank a list of documents by inserting the query and document list into an LLM’s prompt and instructing it to output the reranked document identifiers, though they rely on the strong capability of LLMs, suffer from positional bias and are sensitive to document order in the prompt [49]. Pairwise ranking methods [32] provide the query and a pair of documents to the LLM, which is instructed to generate the identifier of the more relevant document; these pairwise comparisons are then aggregated using efficient sorting algorithms like Heapsort or Bubblesort to produce the final ranking. The Setwise approach [50] is also proposed to compare a set of documents at a time to further improve efficiency.
Rank aggregation has been widely used in many information retrieval tasks [14, 12]. Previous works [31, 15] also use aggregation to form the ranking list from pairwise comparison. We employ Borda count [9] to aggregate different ranking lists in our paper. Borda count assigns a score to each item based on its position in each input ranking and sums these scores to produce the final aggregated ranking. While other rank aggregation methods exist, such as Markov Chain-based methods [12], supervised learning approaches [26], and unsupervised methods like Kemeny rank aggregation [18], many of these are either NP-hard or not specifically designed for list aggregation. By using the simple and computationally efficient Borda count method, we aim to demonstrate the power of combining LLM-based ranking with aggregation, even without resorting to more complex aggregation techniques.
3 Inconsistency of LLM-based ranking
3.1 Inconsistency of Pairwise Comparisons
  1. 1.
  2. 2.
Under the pairwise ranking approach, each LLM query produces a pairwise comparison results on $d_{i},d_{j}$, the result can be $d_{i}\succ d_{j}$ or $d_{j}\succ d_{i}$, where $d$ represents a passage and $\succ$ means "preferred to". While we focus on pairwise comparisons in this work, our analysis of these two types of inconsistency can be easily extended to setwise or listwise ranking schemes. This is because both setwise and listwise comparisons still rely on relative preferences between individual passages, and are therefore fundamentally pairwise in nature. For example, the result of a setwise comparison $d_{i}\succ d_{j},d_{k}$ can be represented as two implicit pairwise comparisons, $d_{i}\succ d_{j}$ and $d_{i}\succ d_{k}$. Therefore, our inconsistency analysis can not only be applied to setwise and listwise ranking, but any ranking scheme that involves relative comparisons between passages.
3.2 Inconsistency Measurement
We represent pairwise comparisons using a tournament graph - a complete graph where each vertex represents a passage and edges represent preferences between passages. In this graph, directed edges $d_{i}\rightarrow d_{j}$ indicate that passage $d_{i}$ is preferred over $d_{j}$, while undirected edges indicate ties. Following PRP-Sort [32], we handle cases of order inconsistency (where swapping passage positions leads to different preferences) by marking them as ties in the graph. While these ties are treated equally in the graph structure, we note that the underlying passages may have different ground truth relevance scores.
To measure transitive inconsistency, we first construct the tournament graph by performing all pairwise comparisons between passages and marking order inconsistent pairs (those that change preference when swapped) as ties. We then count the number of inconsistent triads in the graph, following the method of Kułakowski [19]. An inconsistent triad occurs when three passages form a cycle of strict preferences - for example, when $d_{1}$ is preferred over $d_{2}$, $d_{2}$ is preferred over $d_{3}$, but $d_{3}$ is preferred over $d_{1}$, violating transitivity. The count of such triads serves as our metric for transitive inconsistency. As shown in Table 2, we observe that the frequency of inconsistent triads varies across different LLMs, with larger models generally exhibiting fewer transitive inconsistencies.
@@ 3.3 The Impact of Inconsistency
BM25 Hard List BM25 Hard List
@@
Table 3: NDCG@10 of ranking from different initial orders. Hard list is the inverse order of the ranking result obtained by performing the PRP-Allpair ranking scheme [32]. Note that Heapsort results are largely unaffected, suggesting that this specific hard list is for Bubblesort only.
As shown in Tables 1 and 2, LLM-based rankers can exhibit significant amounts of inconsistency across judgments. Applying sorting-based ranking schemes based on non-transitive pairwise comparisons can produce volatile result rankings that are highly sensitive to the initial order of candidate passages. This can have particularly adverse effects if that initial ordering is a "hard list" for the chosen sorting method, as demonstrated in Table 3. A hard list is an initial order of passages where the high-relevance passages require many comparisons to be moved to the front of the ranking, increasing the likelihood of encountering a transitive inconsistency which blocks the promotion of the passage. A hard list for one sorting algorithm may not be as hard of a list for another, which is demonstrated in Table 3. By aggregating the full ranked lists from multiple sorting algorithms, we can mitigate the worst-case effects of inconsistency while producing more robust final rankings.
@@ 4.1.1 In-Context Learning (ICL)
We design the ICL prompt to utilize the few-shot ability [4] of LLMs to mitigate order inconsistency. The prompt provides the LLM with an example pairwise comparison for both order permutations as shown in Figure 1. This demonstration illustrates that the task is to compare the passages based on their relevance to the query instead of its position in the prompt. As shown in Table 1, using ICL can balance the probability of LLM selecting a passage from either position.
@@ 4.2 LLM Ranking Aggregation
In passage ranking, we can generate multiple proposed ranking lists using different ranking settings. These settings can include different sorting algorithms to build fully ranked lists from pairwise, setwise, or listwise comparisons, as well as other factors like a specific LLM used for the preference query. Given the inconsistency discussed in 3, we know that any configuration of these settings can result in noisy rank results. We cannot assume we know a priori which proposed rank setting is the best, so choosing a single setting that will be affected the least by inconsistency is difficult.
@@
5:     $L_{i}\leftarrow$ Ranking($q$, $D$, $R_{i}$) $\triangleright$ Generate rank list using rank setting $R_{i}$
@@ 4.2.2 Aggregation Across LLMs
Table 4: NDCG@10 of PRP-Sorting with Bubblesort, starting from the BM25 initial order on TREC DL 2019. The experiment shows an ablation study of using ICL and calibration to improve ranking performance by addressing order inconsistency.
DL19 DL20
Setwise 63.25 60.27
Listwise (RankGPT) 69.61 65.49
Pairwise (PRP-Sort) 65.38 60.16
Pairwise+ICL+Calibration (Ours) 71.51 65.10
Table 5: NDCG@10 on TREC DL 2019 and 2020 datasets using Llama-3-8B model. Our proposed method (Pairwise+ICL+Calibration) shows superior performance compared to setwise, listwise, and standard pairwise approaches.
5 Experiments
5.1 Experimental Setup
We utilize test sets from TREC, a standard dataset for information retrieval research. Specifically, we use the top 100 passages retrieved by BM25 [22, 34] for each of the queries associated with the TREC-DL2019 and 2020 test sets [8]. Our results are based on the re-ranking of these 100 passages. The LLM ranking scheme is implemented under the same experimental setting of PRP-Sort [32].
We evaluate our results using two metrics: Normalized Discounted Cumulative Gain (NDCG) and average Kendall tau Distance. NDCG is a standard metric used to evaluate the quality of ranked retrieval results. It accounts for the position of relevant documents in the ranking, assigning higher importance to documents appearing earlier. Kendall tau distance is a metric used to measure the dissimilarity between two rankings. We compare our results against pairwise, setwise, and listwise baselines [32, 50, 38].
@@ 5.2 Addressing Order Inconsistency
We have shown that in-context learning can help to balance the average probability of the two choices given during pairwise ranking in Table 1. In Table 4, we can see that solely using ICL can help improve the ranking performance in most LLMs. Some encoder-decoder models like Flan-UL2 may intrinsically suffer from less positional bias than other decoder-only LLMs [24], so we don’t expect much improvement with using ICL for those models. The calibration addresses order inconsistency by calculating preference probability based on comparison from both positions. The improvement from solely using the calibration is also significant, as shown in Table 4. We also find that using these two methods individually cannot improve ranking performance for Flan-UL2, but combining them can bring improvement compared to the baseline. Furthermore, as demonstrated in Table 5, our approach of combining pairwise ranking with ICL and calibration outperforms several baseline methods across different datasets.
Flan-T5-XXL Llama-3-8B Model Aggregation
TREC DL 2019 Pairwise.BubbleSort 71.05 71.51 73.08
Pairwise.HeapSort 69.76 70.82 72.24
Setwise.BubbleSort 71.07 63.25 64.01
Setwise.HeapSort 70.46 65.14 67.32
Scheme Aggregation 70.78 71.61
TREC DL 2020 Pairwise.BubbleSort 69.96 65.10 70.95
Pairwise.HeapSort 69.61 64.01 69.88
Setwise.BubbleSort 68.66 60.27 61.73
Setwise.HeapSort 68.82 60.68 63.9
Scheme Aggregation 69.74 65.09
Table 6: NDCG@10 of different ranking schemes and model aggregations on TREC DL 2019 and 2020 datasets. The table shows individual performance of Flan-T5-XXL and Llama-3-8B models, as well as their aggregation, across various ranking schemes. The pairwise ranking results in this table both apply ICL and calibration. The initial ranking list is from the BM25 order.